Three individuals form a partnership and agree to divide the profits equally. X invests $9,000, Y invests $7,000, Z invests $4,000. If the profits are $4,800, how much less does X receive than if the profits were divided in proportion to the amount invested?
Answer by mananth(16946)

Total investment by all three = $20000
x investment = $9000
ratio of investment = 9000/20000= 9/20
y investment = $7000
ratio of investment = 7000/20000 = 7 / 20
z= investment =$4000
ratio of invetment = 4000/20000=1/5
PROFIT = $4800
the profit has to be shared in the ratio of their investment
A= 9/20 * 4800= $ 2160
Y = 7/20 * 4800 = $1680
Z= 1/5 * 4800 = $960